OpenStudy (anonymous):

A stadium has 52,000 seats. Seats sell for $42 in section A, $36 in section B, and $30 in section C. The number of seats in section A equals the total number of seats in sections B and C. Suppose the stadium takes in $1,960,200 from each sold out event. How many seats does each section hold?

OpenStudy (anonymous):

42A+36B+30C=1,960,200 and A=B+C and A+B+C=52,000

OpenStudy (anonymous):

That might help start